log2X+log2(x+4)-log2(x-2)=4
Answer by lwsshak3(11628)

log2X+log2(x+4)-log2(x-2)=4
log2[x(x+4)/(x-2)]=log2(16)
x(x+4)/(x-2)=16
x^2+4x=16x-32
x^2-12x+32=0
(x-8)(x-4)=0
x=8
or
x=4